OpenStudy (anonymous):

what are positive integers for n such that 36+2∧n is a perfect square

OpenStudy (anonymous):

is it\[36+2^n\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

one answer is n=6

OpenStudy (anonymous):

\[2^5+2^n\]

OpenStudy (anonymous):

hmm are you sure? Because 2⁵ is 32

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait your ? is 36 my bad

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

\[k^2=6^2+2^n\] \[k^2-6^2=2^n\] \[(k-6)(k+6)=2^n\]k>6 @dplanc

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

surely k=10 means n=6